solve
5 1/2x - 7 1/3 = 9 1/2

Answer by stanbon(75887)

5 1/2x - 7 1/3 = 9 1/2
(11/2)x - (22/3) = (19/2)
Multiply thru by 6 to get rid of the denominators:
33x - 44 = 57
33x = 101
x = 101/33
================
Cheers,
Stan H.